OpenStudy (anonymous):

no \[\pm \frac{ \frac{ 1 }{ 2 } }{ \frac{ \sqrt{3} }{ 2 } }=\pm \frac{ 1 }{ \sqrt{3} }=\pm \frac{ \sqrt{3} }{ 3 } = \pm \frac{ 1.732 }{ 3 }=?\]

OpenStudy (anonymous):

\[\sin x=\frac{ 1 }{ 2 }\] \[\sin ^2x+\cos ^2x=1,\cos ^2x=\frac{ 3 }{ 4 }\] \[\cos x=\pm \frac{ \sqrt{3} }{ 2 }\] \[\tan x=\pm ?\]

OpenStudy (abb0t):

When you have a square root, for example, you have \(\sf \sqrt{3}\) what that actully is, is a exponent, so you have \(\sf 3^{\color{red}{{\frac{1}{2}}}}\)

OpenStudy (abb0t):

Oh, nevermind, ignore that. Memorize, SOH CAH TOA SOH = sine = opposite over hypotenuse TOA = tan = opposite over adjacent etc.,